Find the equation of the line passing through the point (2,3) and at a distance of 3 units from the origin.

Grade:

1 Answers

Abhishek Singh
9 years ago
let eqn. of line be  y-3=m(x-2), where m is slope of line 
 dist. from origin = 3
 I0-3-m(0-2)I/(1+m2)1/2=3
solving we get , m=0 or -12/5
 
lines are y=0 or 5y+12x=39